### v4

@@ Abstract
keywords:
[…]
1 Introduction
Lane detection plays an important role in Advanced Driver Assistance Systems (ADAS), providing essential support for lane keeping, autonomous navigation, and lane departure warnings, which help reduce the risk of lane-deviation accidents. As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving.
Under good weather and clear lane markings, current lane detection algorithms perform well. However, in complex road environments and poor weather, such as fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging. In these conditions, existing algorithms often struggle and may fail, leading to an increased risk of traffic accidents with more severe consequences[1]. Ensuring safe driving in these conditions calls for the development of high-precision, highly robust lane detection methods optimized for foggy environments. As shown in Fig. 1, the in-vehicle camera captures real-time road images, which are preprocessed by edge devices and then sent to the lane detection module. Detected lane information is fused with data from other sensors, such as radar and GPS, along with real-time traffic data from the cloud. This information is processed in the vehicle’s computing unit, allowing the system to make adaptive driving decisions such as accelerating, braking, or steering[2, 3] to ensure safety in challenging environments like fog.
Although end-to-end autonomous driving technologies have made significant progress at Level 3 (L3) and higher, most current vehicles, particularly commercial and urban vehicles, are still at Level 0 (L0), Level 1 (L1), and Level 2 (L2). In these lower levels of autonomous driving, lane detection remains a crucial technology, especially in complex and adverse road environments. Most existing vehicles rely on Advanced Driver Assistance Systems (ADAS) to enhance safety in complex scenarios, especially in active warning, lane-keeping, and other functions[4]. Therefore, this research not only focuses on end-to-end autonomous driving technologies but also aims to improve lane detection performance in complex environments, providing stronger safety support and driving assistance for L0, L1, and L2 vehicles in adverse conditions.
@@ 1 Introduction
Earlier lane detection algorithms were mainly based on traditional computer vision methods, typically including three steps: image preprocessing, feature extraction, and lane model fitting. Son et al.[6] first detected the vanishing point to adaptively select the ROI, then extracted lane markings based on color features, followed by clustering to find the main lane. Hou et al.[7] used the Canny operator to extract lane edges, filtered out noise with abnormal directions, and applied a progressive probabilistic Hough transform for line detection, filtering out unwanted lines with vanishing point constraints. These feature-based methods are simple, efficient, and suitable for real-time applications, but they also have clear limitations, such as requiring manual parameter tuning and high dependence on clear lane edges. When lane markings are worn, obscured, or subject to environmental changes, accurate detection becomes challenging, and robustness is limited. Improving and optimizing these methods remains necessary to enhance their robustness and adaptability for practical use.
Owing to the rapid development of deep learning, it has been widely applied in various autonomous driving tasks. For example, in traffic object detection, Wang and Zhang[8] proposed CRRFNet, an adaptive method that combines camera and radar fusion, improving detection accuracy and robustness in noisy or abnormal scenes. In accident detection and classification, Yang et al.[9] developed a deep learning model that utilizes multi-vehicle trajectory data for detecting and classifying accidents on freeways, achieving up to 100% detection accuracy. For 3D object detection under adverse weather, Qi et al.[10] introduced a novel geometric information constraint network (GIC-Net), enhancing the performance of 3D object detection by incorporating geometric location and line feature constraints, demonstrating superior results under challenging weather conditions. In addition to these applications, lane detection has also seen significant advancements, with deep learning-based approaches becoming mainstream due to their strong feature extraction capabilities and generalization. Compared to traditional methods, deep learning provides higher robustness, better adapting to complex road environments and effectively handling cases with missing or occluded lane markings. LaneNet[11] approaches lane detection as a combination of semantic segmentation and post-processing, first generating a semantic segmentation map of the lanes and then using density clustering to identify each lane instance. However, because of the large number of pixels generated during segmentation and the inability of clustering to leverage GPU acceleration, its real-time performance remains limited. LaneAF[12], building on the semantic segmentation map, introduced a clustering algorithm based on horizontal and vertical vector fields to accelerate post-processing and used DLA[13] as the backbone network to integrate multi-level semantic and spatial information, enhancing feature representation. SCNN[14] leveraged the elongated structure of lane lines by replacing traditional layer-wise convolution with slice convolution, enabling cross-row and cross-column information transfer, which makes it well-suited to line-like lane structures. However, its large parameter size prevents real-time performance. Overall, current networks, whether for semantic or instance segmentation, mostly rely on pixel-by-pixel classification, though lane detection does not necessarily require dense pixel segmentation. To address this, Qin et al.[15] reframed lane detection as a row-based segmentation problem using global features, significantly reducing computational cost and improving detection speed by selecting specific rows and classifying each pixel in the row. This method also leverages the large receptive field of global features, enabling it to handle complex scenarios with missing or occluded lanes.
Apart from segmentation, lane detection also includes detection-based methods, many of which borrow the Anchor-based Methods from object detection[16, 17, 18, 19]. However, due to the elongated shape of lane lines, traditional bounding boxes are not ideal. In lane detection, anchors are often a series of preset starting points and straight lines with different slopes, represented as point sets, with the network predicting the relative distance between the lane line and anchor to optimize line shape and position. PointLaneNet[20] first represents lane lines as a series of points, taking the intersections of evenly spaced horizontal lines as coordinates $\{x_{1},x_{2},\ldots,x_{n}\}$ to describe the lane. Like other YOLO-based methods[16, 17], PointLaneNet divides the image into multiple grids, predicts lane position and classification at each grid center, and removes duplicate predictions through post-processing. However, this grid-based prediction approach has limitations for elongated lanes, reducing detection accuracy. LaneATT[21] addresses this issue by noting the high correlation between lane positions and introducing global features to infer lane position, especially in cases of occlusion or missing lane markers. LaneATT[21] introduced an anchor-based attention mechanism that significantly improves detection speed and accuracy by utilizing global information. CLRNet[22] demonstrated the complementary roles of low- and high-dimensional features in lane detection, proposing Cross Layer Refinement (CLR) to dynamically adjust anchor points through network learning, removing the limitations of fixed anchors to optimize lane detection further. FLAMNet[23] proposes an adaptive line anchor framework integrated with a dual-branch network design to improve the accuracy and adaptability of capturing intricate lane geometries. By leveraging the Decomposed Self-Attention (DSA) mechanism alongside a Horizontal Feature Fusion Module (HFFM), the model enhances the perceptual range of line anchors in lane detection tasks. This approach strengthens robustness against occlusions and incomplete lane markings in challenging environments.
In addition to these methods, trajectory data has also been explored for lane identification. Zhao and Roncoli[24] proposed a constrained spectral clustering method to identify lane information from high-resolution trajectory data, significantly improving results even in the presence of varying lane numbers or lane-changing maneuvers. This approach demonstrates the potential of leveraging trajectory data for lane detection, especially in more dynamic environments.
@@
Compared to other detection and segmentation tasks, research on downstream tasks in adverse weather conditions, such as fog, is relatively limited. A conventional solution often applies standard image enhancement techniques, such as classical dehazing methods developed for fog removal and visual quality improvement. Nevertheless, enhanced visual clarity doesn’t always correlate with superior detection outcomes. Some methods combine image enhancement with detection to reduce adverse weather interference. Huang et al.[25] proposed DSNet, incorporating a detection subnet and recovery subnet, where multi-task learning combines dehazing and detection to improve understanding of foggy scenes, resulting in better detection than a standard detection network. Additionally, some methods simulate foggy images from existing datasets to optimize downstream tasks. For example, Nie et al.[26] used the Monodepth2 model and atmospheric scattering model to convert clear lane images from the CULane dataset into foggy lane images, achieving synthetic data augmentation. Testing with SCNN[14] showed that this augmentation strategy improved recognition rates from 76.65% to 86.65% under light fog conditions.
Another approach employs domain adaptation techniques to address the style gap between foggy scenes and clear-weather images[27, 28]. Hnewa et al.[29] assumed that there is domain shift between images captured in normal and adverse weather conditions, proposing a multi-scale domain-adaptive YOLO that supports domain adaptation at different levels in the feature extraction stage. Li et al.[30] proposed a novel transfer learning-based framework for fog-adaptive object detection, which integrates dual-level adaptation (image-level and object-level) to mitigate domain shifts caused by foggy conditions. Notably, the method only requires labeled clear-weather images and unlabeled foggy data for training, eliminating the dependency on annotated foggy datasets.
@@
Traditional dehazing methods, which are based on statistical or physical principles, often suffer from artifacts such as image darkening or brightening. In contrast, end-to-end deep learning methods using Convolutional Neural Networks (CNNs) can capture more complex features but come with higher computational and memory demands. These methods are also highly dependent on diverse training datasets. However, acquiring real paired images under foggy conditions is challenging, and as a result, many existing datasets rely on synthetic images, which can negatively affect algorithm performance in real-world scenarios.
@@ 2 Related Work
2.1 Segmentation-based Lane Detection
Segmentation-based methods mainly utilize semantic segmentation or instance segmentation techniques to distinguish lane lines from other objects or backgrounds in the image, transforming lane detection into a pixel-level classification problem. LaneNet[11] employs a two-branch structure of split branches and embedded branches, the split branches output a binary lane mask to recognize lane lines, and the embedded branches assign a unique embedding to each pixel to realize the distinction of lane lines. SCNN[14] extends the traditional layer-by-layer convolution operation with a slice-by-slice convolution operation, which passes the information horizontally and vertically between the pixels within a layer, thus enhancing the continuity of lane line recognition. RESA[32] introduces a spatial attention module that ensures robust lane prediction by cyclically shifting the feature map horizontally and vertically so that each pixel can acquire global information. Although these segmentation-based methods possess high accuracy and robustness, the models are bulky and slow to process due to high consumption of computational resources.
2.2 Curve-based Lane Detection
Curve-based methods transform the lane detection into a regression task by predicting polynomial coefficients that fit lane lines to polynomial equations. PolyLaneNet[33] introduces an end-to-end convolutional neural architecture that maps from images to polynomials representing each lane marker in the image, as well as domain lane polynomials and confidence scores for each lane. LSTR[34] can also directly output the parameters of the lane shape model, has developed a Transformer-based network that takes into account non-local interactions, enabling it to capture the narrow structure of the lane lines and global context information. DBNet[35] introduces NURBS curves and utilizes their geometric semantic properties to achieve local and global optimization, which improves the robustness of lane line detection and model interpretation. However, polynomial fitting is sensitive to parameter selection, and minor changes in parameters can result in significant variations in fitting results, leading to unstable lane line detection.
2.3 Anchor-based Lane Detection
@@
2.4 Row-wise-based Lane Detection
@@ 3 Data Constrution
3.1 FoggyLane
In the field of intelligent connected vehicle ADAS systems, existing lane detection algorithms have shown good performance under favorable weather conditions. However, their performance remains significantly inadequate under adverse weather conditions, such as fog. Given that fog is a high-risk scenario for traffic accidents, the limited visibility drastically reduces the driver’s reaction time and decision-making ability, making accidents more likely. However, current benchmark datasets for lane detection, including CULane[14], Tusimple[31], and LLAMAS[40], do not include data from foggy scenarios. To address this gap, we have developed a lane detection dataset, FoggyLane, specifically designed for foggy conditions. Given the randomness of foggy weather, data collection was carried out using a monocular forward-facing camera at different times and locations. This includes the morning of October 26, 2023, in Suzhou, Jiangsu Province, China, and two sessions in Urumqi, Xinjiang Uygur Autonomous Region, China: the morning of December 9, 2023, and the night of January 2, 2024. In these selected representative traffic scenarios, we sampled one frame every frames, resulting in a total of 1,086 RGB images: 444 images (resolution: $1920\times 1080$), 102 images (resolution: $2560\times 1920$), and 540 images (resolution: $2304\times 1296$). To simplify operations and enhance usability during training, we resized all images to a uniform resolution of $1640\times 590$.
However, since the duration of foggy weather conditions is limited, the number of collected foggy weather images is restricted. To improve the model’s performance, we extended our dataset by incorporating additional foggy driving scenes. Specifically, we downloaded two foggy driving videos from YouTube[41], extracting road-driving segments at 20-frame intervals, obtaining 44 and 293 frames, respectively. This added a total of 337 additional RGB images to the dataset. The original resolution of these images was $1920\times 1080$, and they were resized to $1640\times 590$ for consistency.
As a result, we have constructed FoggyLane, a foggy weather lane detection dataset consisting of 1,423 images. FoggyLane includes six types of road scenes, as shown in Fig. 4, covering Normal, Arrow, Crowd, Curve, Night and Crossroad. We split the dataset into training and testing sets at a ratio of 2:1. To ensure comprehensive learning of scene characteristics, each type of scene is evenly distributed between the training and testing sets to maintain balance.
Table 1: Comparison Between Our Dataset and Existing Lane Detection Datasets
@@ 3.1 FoggyLane
Figure 4: The different road types of FoggyLane images examples.
Meanwhile, We compared FoggyLane with existing lane detection datasets. As shown in Table 1, the Caltech-Lanes and Tusimple datasets only cover single traffic scenarios. The LLAMAS [40] dataset focuses on highway scenarios, but its annotations are generated from high-precision maps rather than manually annotated. While the ApolloScape [43] and BDD100k [44] datasets contain multiple traffic scenarios, they lack instance-level annotations, and ApolloScape [43] does not provide publicly available test set annotations. In comparison, although CULane [14] and CurveLanes [45] include a wide range of traffic scenes with comprehensive annotations, they still lack foggy weather scenarios. For lane detection tasks, models trained only on clear weather conditions struggle to handle the challenges posed by fog, such as blurred lane boundaries, faded colors, and reduced image quality. To address these limitations, we developed FoggyLane, a comprehensive foggy weather lane detection dataset that includes 1,423 images featuring various challenging scenarios such as foggy roads during the day, curves, crossroads, and night fog. The dataset is meticulously annotated using the LabelMe tool, providing instance-level annotations, and it is available in Tusimple [31] and CULane [14] formats for broader accessibility.
@@
Figure 5: Networks of foggy image generation. $\otimes$ represents element-wise multiplication.
Figure 6: The FoggyCULane dataset examples. (a) is the original CULane dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
3.2 FoggyCULane and FoggyTusimple
@@
The method for generating foggy images is inspired by the model [26], and the complete generation process is shown in Fig. 5. To be specific, the atmospheric scattering model can effectively utilize prior information to analyze the degradation mechanisms of image quality in foggy conditions. It models foggy images as a combination of scattered light and direct light based on optical principles, ultimately synthesizing a clear image into a foggy one. The mathematical modeling equation defined by Koschmieder[47] is represented as:
$I(x)=J(x)t(x)+A(1-t(x))$ (1)
where $x$ indicates a certain pixel of the image, $J(x)$ refers to the original clear image of the object and $I(x)$ refers to the foggy image. $A$ is the atmospheric light value at infinity and $t(x)$ can be further defined as:
$t(x)=e^{-\beta d(x)}$ (2)
where $\beta$ is the extinction coefficient, which is closely related to the concentration, size, type and distribution of particulate matter. $d(x)$ represents the distance between the object and the viewer. As this distance increases, the transmittance decreases, resulting in thicker fog.
Figure 7: The FoggyTuSimple dataset examples. (a) is the original TuSimple dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
To calculate the corresponding $d(x)$, we use the S2R-DepthNet[48] that has good generalization performance across various applications. The S2R-DepthNet consists of three modules that decouple the structural and textural information of the input image. Then they suppress information irrelevant to depth and use structural information for depth prediction. The Structure Extraction Module (STE) decomposes the image to obtain domain-invariant structure information. The Depth-specific Attention (DSA) predicts an attention map that helps to suppress depth-irrelevant structures by leveraging high-level semantic information extracted from the raw image. After passing through the DSA Module and STE Module, as shown in Fig. 5, the attention map and the structure map are multiplied to obtain a concise and comprehensive depth-specific structure map. The Depth Prediction (DP) is responsible for depth estimation.
Figure 8: Overview of the proposed network.
According to the transmittance formula Eq. (2), we use the dark channel prior[49] to calculate the average atmospheric light value $A$. In the non-sky area of an outdoor clear image, there is usually at least one very small or close to zero value in the RGB image for each pixel, which is called the dark channel of that pixel. Therefore, for any image $I$, the mathematical formula for its corresponding dark channel image $I_{\text{dark }}(x)$ can be defined as:
$I_{\text{dark }}(x)=\min_{c\in R,G,B}\left(\min_{y\in\Omega(x)}I_{c}(y)\right)$ (3)
where $I_{\text{c}}(y)$ is the value of the c-th channel of image $I$ at pixel point $y$, and $\omega(x)$ is a small region around pixel point $x$. In the dark channel, find the top 0.1% brightest pixels and take the maximum value of their RGB channels as an approximation of the average atmospheric light value $A$. After obtaining the transmittance and the average atmospheric light value, the images of light fog, medium fog, and heavy fog can be obtained by fine-tuning the extinction coefficient $\beta$. The original images and the foggy images generated under different $\beta$ values are displayed in Figs. 6 and 7.
4 Methods
4.1 Overall structure

### v5

@@ Abstract
Index Terms:
ADAS, intelligent vehicle, foggy lane detection dataset, foggy lane detection, global attention, edge enhancement
I Introduction
Lane detection plays an important role in Advanced Driver Assistance Systems (ADAS), providing essential support for lane keeping, autonomous navigation, and lane departure warnings, which help reduce the risk of lane-deviation accidents. As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving[1].
Under good weather and clear lane markings, current lane detection algorithms perform well. However, in complex road environments and poor weather, such as fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging. In these conditions, existing algorithms often struggle and may fail, leading to an increased risk of traffic accidents with more severe consequences. Ensuring safe driving in these conditions calls for the development of high-precision, highly robust lane detection methods optimized for foggy environments. As shown in Fig. 1, the in-vehicle camera captures real-time road images, which are preprocessed by edge devices and then sent to the lane detection module. Detected lane information is fused with data from other sensors, such as radar and GPS, along with real-time traffic data from the cloud. This information is processed in the vehicle’s computing unit, allowing the system to make adaptive driving decisions—such as accelerating, braking, or steering—to ensure safety in challenging environments like fog.
@@ I Introduction
Earlier lane detection algorithms mainly relied on traditional computer vision methods, involving image preprocessing, feature extraction, and lane model fitting. Son et al.[3] first detected the vanishing point to select the ROI adaptively, then extracted lane markings based on color features and performed clustering. Hou et al.[4] used the Canny operator for edge detection and applied a progressive probabilistic Hough transform to detect lines. While these feature-based methods are efficient and suitable for real-time applications, they require manual parameter tuning and struggle in challenging conditions such as worn or occluded lane markings.
With deep learning’s success in image tasks, it has become the mainstream approach for lane detection due to its strong feature extraction capabilities and better generalization. Deep learning offers higher robustness, adapting well to complex road environments and handling missing or occluded lane markings[5, 6, 7, 8, 9, 10]. However, many networks rely on pixel-by-pixel classification, whereas lane detection does not require dense pixel segmentation. Qin et al.[11] addressed this by reframing lane detection as a row-based segmentation problem, reducing computational cost and improving detection speed.
In addition to segmentation methods, lane detection also includes detection-based approaches, many borrowing Anchor-based methods from object detection[12, 13, 14, 15]. Due to the elongated shape of lane lines, traditional bounding boxes are unsuitable. In lane detection, anchors are often preset starting points and lines with different slopes, with the network predicting the relative distance between the lane and the anchor to optimize shape and position[16, 17, 18, 19, 20, 21, 22]. These innovations improve accuracy and efficiency in complex or occluded scenarios.
@@
Research on downstream tasks in adverse weather conditions[23], such as fog, is limited compared to other detection and segmentation tasks. Traditional solutions often use image enhancement techniques, like dehazing, to improve visual clarity, but this doesn’t always lead to better detection results. Some methods combine image enhancement with detection to reduce weather interference[24], while others simulate foggy images to augment datasets. For example, Nie et al.[25] generated foggy images from clear lane images in the CULane dataset, achieving synthetic data augmentation that improved recognition rates from 76.65% to 86.65% under light fog.
Domain adaptation techniques also address the style gap between foggy and clear-weather images[26, 27]. Hnewa et al.[28] proposed a multi-scale domain-adaptive YOLO model, while Li et al.[29] introduced a transfer learning framework for fog-adaptive object detection using labeled clear-weather and unlabeled foggy images, avoiding the need for annotated foggy datasets.
@@
Traditional dehazing methods suffer from artifacts, whereas deep learning methods using CNNs capture more complex features but require high computational resources and diverse training datasets. However, obtaining real paired images under foggy conditions is challenging, and synthetic images often lead to reduced performance in real-world scenarios.
@@ II Related Work
II-A Segmentation-based Lane Detection
Segmentation-based methods mainly utilize semantic segmentation or instance segmentation techniques to distinguish lane lines from other objects or backgrounds in the image, transforming lane detection into a pixel-level classification problem. LaneNet[5] employs a two-branch structure of split branches and embedded branches, the split branches output a binary lane mask to recognize lane lines, and the embedded branches assign a unique embedding to each pixel to realize the distinction of lane lines. SCNN[8] extends the traditional layer-by-layer convolution operation with a slice-by-slice convolution operation, which passes the information horizontally and vertically between the pixels within a layer, thus enhancing the continuity of lane line recognition. RESA[31] introduces a spatial attention module that ensures robust lane prediction by cyclically shifting the feature map horizontally and vertically so that each pixel can acquire global information. Although these segmentation-based methods possess high accuracy and robustness, the models are bulky and slow to process due to high consumption of computational resources. HW-Transformer[32] presents an innovative lane detection network that utilizes row and column multi-head self-attention to expand the visual range around the lane and enable global information communication through intersecting features, alongside a self-attention knowledge distillation (SAKD) method to enhance performance and semantic feature learning.
II-B Curve-based Lane Detection
Curve-based methods transform the lane detection into a regression task by predicting polynomial coefficients that fit lane lines to polynomial equations. PolyLaneNet[33] introduces an end-to-end convolutional neural architecture that maps from images to polynomials representing each lane marker in the image, as well as domain lane polynomials and confidence scores for each lane. LSTR[34] can also directly output the parameters of the lane shape model, has developed a Transformer-based network that takes into account non-local interactions, enabling it to capture the narrow structure of the lane lines and global context information. DBNet[35] introduces NURBS curves and utilizes their geometric semantic properties to achieve local and global optimization, which improves the robustness of lane line detection and model interpretation. However, polynomial fitting is sensitive to parameter selection, and minor changes in parameters can result in significant variations in fitting results, leading to unstable lane line detection. PGA-Net[36] introduces an end-to-end framework for lane detection, utilizing cubic polynomial functions and a transformer-based DETR model to simultaneously predict lane shape parameters and incorporate global road context, with the addition of Mean Curvature Loss to enhance the accuracy of curve predictions.
II-C Anchor-based Lane Detection
@@
II-D Row-wise-based Lane Detection
@@ III Data Constrution
III-A FoggyLane
Existing lane detection algorithms for intelligent connected vehicle ADAS systems perform well under favorable weather conditions, but their performance significantly deteriorates in adverse conditions like fog, which reduces visibility and increases the likelihood of traffic accidents. Current benchmark datasets, including CULane[8], Tusimple[30], and LLAMAS[41], lack foggy weather data. To address this, we developed FoggyLane, a dataset specifically designed for foggy conditions. Data was collected using a monocular forward-facing camera at various times and locations, including Suzhou and Urumqi, China, between October 2023 and January 2024. The dataset includes 1,086 images at various resolutions, which were resized to $1640\times 590$ for consistency.
Since foggy conditions are rare, we extended the dataset by incorporating foggy driving scenes from two YouTube videos[42], resulting in an additional 337 images. The final FoggyLane dataset contains 1,423 images, covering six road scenes: Normal, Arrow, Crowd, Curve, Night, and Crossroad. We split the dataset into training and testing sets at a 2:1 ratio, ensuring balanced distribution across scene types.
Comparing FoggyLane with existing datasets, such as Caltech-Lanes[43] and Tusimple[30], which focus on single traffic scenarios, and LLAMAS[41], which uses high-precision map annotations, we found that these datasets do not cover foggy weather scenarios. While ApolloScape [44] and BDD100k [45] include multiple traffic scenarios, they lack instance-level annotations. Similarly, although CULane [8] and CurveLanes [46] cover a wide range of traffic scenes, they do not include foggy conditions. Models trained on datasets limited to clear weather struggle in fog, where lane boundaries are blurred and image quality reduced. To address this gap, FoggyLane provides a comprehensive solution for foggy weather lane detection with instance-level annotations, available in both Tusimple [30] and CULane [8] formats for broader accessibility.
Figure 4: The different road types of FoggyLane images examples.
TABLE I: Comparison Between Our Dataset and Existing Lane Detection Datasets
@@ III-A FoggyLane
Figure 5: The FoggyCULane dataset examples.
III-B FoggyCULane and FoggyTusimple
@@
The method for generating foggy images is inspired by the model [25], and the specific generation method is provided in the Supplementary Material for Review. As shown in Figure 5, it demonstrates the generation case of FoggyCULane.
IV Methods
IV-A Overall structure